OpenStudy (anonymous):

Select the system below which meets the following conditions: The slope-intercept forms are identical, so the graphs are identical. Both equations share all their ordered pairs as solutions. A. y + 3x = 5 y = 2x + 5 B. y = 4x - 1 y = 3x - 2 C. y = 2x - 4 y/2 + 2 = x D. None of the above

OpenStudy (anonymous):

or rather C

OpenStudy (anonymous):

\[\frac{y}{2}+2=x\] \[\frac{y}{2}=x-2\] \[y=2x-4\]